### v1

# Batch Acquisition Function Evaluations and Decouple Optimizer Updates for Faster Bayesian Optimization

Abstract
Bayesian optimization (BO) efficiently finds high-performing parameters by maximizing an acquisition function, which models the promise of parameters. A major computational bottleneck arises in acquisition function optimization, where multi-start optimization (MSO) with quasi-Newton (QN) methods is required due to the non-convexity of the acquisition function. BoTorch, a widely used BO library, currently optimizes the summed acquisition function over multiple points, leading to the speedup of MSO owing to PyTorch batching. Nevertheless, this paper empirically demonstrates the suboptimality of this approach in terms of off-diagonal approximation errors in the inverse Hessian of a QN method, slowing down its convergence. To address this problem, we propose to decouple QN updates using a coroutine while batching the acquisition function calls. Our approach not only yields the theoretically identical convergence to the sequential MSO but also drastically reduces the wall-clock time compared to the previous approaches.
Codehttps://github.com/Kaichi-Irie/aaai2026-anon-exp/tree/submission
1 Introduction
Bayesian optimization (BO) is a prevalent method to reduce trial-and-error iterations in many applications with an expensive objective such as materials discovery (xue2016accelerated; li2017rapid; vahid2018new), hyperparameter optimization (HPO) of machine learning models (loshchilov2016cma; chen2018bayesian; feurer2019automated), and drug discovery (schneider2020rethinking). Since the acquisition function optimization plays a central role in an efficient search, multi-start optimization (MSO) by a quasi-Newton (QN) method, which is essential to enhance the solution quality, is employed by BoTorch (balandat2020botorch) and Optuna (akiba2019optuna).
Although MSO is widely regarded as the gold standard, its overhead can be non-negligible when each optimization is run sequentially (denoted Seq. Opt.), as it multiplies acquisition-function evaluations that each entail an expensive Gaussian-process regressor call. BoTorch optimizes the sum of acquisition-function values across multiple restarts  [^1] See “Multiple Random Restarts” in the BoTorch documentation. URL: https://botorch.org/docs/v0.15.0/optimization/#multiple-random-restarts. to mitigate this cost. This formulation enables PyTorch to batch acquisition function evaluations, reducing overall overhead while preserving the theoretically exact per-point gradients.
Despite these advantages, this formulation substantially slows down the convergence of QN methods based on our experiments. We empirically show that the slowdown is explained by the approximation error in cross derivatives or off-diagonal blocks, which we call off-diagonal artifacts, between dimensions that belong to independent points. Off-diagonal artifacts distort search directions, leading to slower convergence.
As a remedy, we propose Decoupling QN updates per restart while Batching acquisition function Evaluations (D-BE). By contrast, we refer to the BoTorch practice, which Couples QN updates while using Batched Evaluations, as C-BE. To our knowledge, there is no practical block-structure-aware, bound-constrained, which is ubiquitous in BO, QN algorithm that removes these off-diagonal artifacts. D-BE sidesteps this difficulty by keeping per-point states independently while retaining batched evaluations, by combining a coroutine and batching—no new solver is required. In experiments, D-BE eliminates off-diagonal artifacts while leveraging hardware throughput and preserving solution quality, yielding up to $1.5\times$ and $1.1\times$ wall-clock speedups over Seq. Opt. and C-BE, respectively.
In summary, our contributions are as follows:
  1. 1.
    Present a pitfall of C-BE that slows the convergence, focusing on off-diagonal artifacts,
  2. 2.
    Propose D-BE achieved by a novel combination of a coroutine and batching, which avoids the pitfall without degrading solution quality, and
  3. 3.
    Demonstrate that D-BE is faster than Seq. Opt. and C-BE in experiments.
Figure 1: Contour maps of the inverse Hessian (Left) and the inverse Hessians approximated by L-BFGS-B with Seq. Opt. (Center) and C-BE (Right) on the Rosenbrock function, evaluated near the constrained minimizer ($B=3,D=5,\mathbf{x}\in[0,3]^{D}$). Each figure has $15\times 15$ tiles, and the $(i,j)$-th tile corresponds to the colormap for the $(i,j)$-th element of the (approximated) inverse Hessian. Blue and yellow represent lower and higher values, respectively. Each subtitle reports $e_{\mathrm{rel}}(H)=\|H-H_{\mathrm{true}}\|_{F}/\|H_{\mathrm{true}}\|_{F}$. Left: The true inverse Hessian exhibits zero at off-diagonal blocks. Center: The approximated inverse Hessian by Seq. Opt.. Off-diagonal blocks show zero everywhere. Right: The approximated inverse Hessian by C-BE. Off-diagonal blocks are dense because C-BE does not allow QN methods to be aware of the zero off-diagonal block nature.
2 Background
Bayesian Optimization
Given a function $f(\mathbf{x}):\mathbb{R}^{D}\rightarrow\mathbb{R}$ to be maximized, Bayesian optimization iteratively suggests a parameter vector $\mathbf{x}\in\mathbb{R}^{D}$ by maximizing an acquisition function $\alpha:\mathbb{R}^{D}\rightarrow\mathbb{R}$. When the acquisition function is defined on a continuous space, QN methods are widely used for the optimization owing to their fast convergence. Even if the analytical expression of gradients $\nabla\alpha(\mathbf{x})$ is cumbersome to derive, PyTorch enables us to compute the gradients numerically using automatic differentiation (AD) (ament2023unexpected; daulton2020differentiable). Although QN methods achieve fast convergence, the non-convexity of acquisition functions necessitates MSO. Namely, the acquisition function is optimized $B$ times independently to enhance the solution quality. The goal of this paper is to propose a method to speed up the $B$ times of independent optimizations without degrading the solution quality.
Quasi-Newton (QN) Methods
QN methods, such as BFGS, L-BFGS, and L-BFGS-B, optimize a function using first-order gradients together with an approximation to the inverse Hessian, i.e., the matrix of second derivatives (nocedal1980updating; liu1989limited; byrd1994representations; byrd1995limited). In particular, L-BFGS-B is widely used for optimizing acquisition functions because of its fast convergence and ability to handle box constraints.
Multi-Start Optimization: Sequential vs. Batching
Two schemes have been considered for MSO so far, (1) Sequential Optimization (Seq. Opt.), which performs an independent optimization from each starting point sequentially (see Algorithm 2 in Appendix A for more details), and (2) Coupled Updates with Batched Evaluation (C-BE). More concretely, C-BE optimizes the summation of acquisition function values from independent restarts. Namely, C-BE optimizes the following function with QN methods:
$\alpha_{\mathrm{sum}}(\mathbf{X})\coloneqq\sum_{b=1}^{B}\alpha\big(\mathbf{x}^{(b)}\big),$ (1)
where $\mathbf{x}^{(b)}\in\mathbb{R}^{D}$ is a point in the $b$-th optimization and $\mathbf{X}=\{\mathbf{x}^{(b)}\}_{b=1}^{B}\in\mathbb{R}^{B\times D}$ is a matrix with each point from the $b$-th optimization at the $b$-th column. Since $\alpha_{\mathrm{sum}}$ is additively separable, the gradient of $\alpha_{\mathrm{sum}}$ with respect to $\mathbf{x}^{(b)}$ matches that of $\alpha(x^{(b)})$, enabling a batched evaluation to get the acquisition function values and their gradients at multiple points simultaneously. In a nutshell, C-BE optimizes a $BD$-dimensional (additively separable) function instead of solving a $D$-dimensional function $B$ times independently. BoTorch uses this technique to accelerate MSO.
3 Pitfall of C-BE: Off-Diagonal Artifacts
Although C-BE has been used to accelerate Bayesian optimization, it can in fact slow down each BO iteration or degrade the solution quality when it is used with QN methods.
When using first-order (gradient-based) methods, each block of $\nabla_{\mathbf{X}}\alpha_{\mathrm{sum}}(\mathbf{X})$ equals the per-restart gradient $\nabla\alpha(\mathbf{x}^{(b)})$, making the convergence of C-BE comparable to that of Seq. Opt. By contrast, QN methods approximate the inverse Hessian, and the accuracy of this approximation strongly affects the convergence rate. Importantly, the Hessian (and its inverse) of the summed acquisition over multiple restarts has zero cross-partial derivatives between variables associated with different restarts. More precisely, the true Hessian of $\alpha_{\mathrm{sum}}$ is block-diagonal:
$H\coloneqq\nabla_{\mathbf{X}}^{2}\alpha_{\mathrm{sum}}(\mathbf{X})=\begin{bmatrix}\nabla^{2}\alpha(\mathbf{x}^{(1)})&&\mathbf{0}\\ &\ddots&\\ \mathbf{0}&&\nabla^{2}\alpha(\mathbf{x}^{(B)})\end{bmatrix}$. (2)
However, structure-oblivious QN updates in the $BD$-dimensional space typically maintain a dense inverse-Hessian approximation, injecting non-zero values into off-diagonal blocks and distorting search directions. We refer to this phenomenon as off-diagonal artifacts. Figure 1 visualizes the inverse Hessian and its L-BFGS-B approximation (memory size $m=10$) under Seq. Opt. and C-BE. The figure intuitively shows C-BE induces off-diagonal artifacts.
The next question is: “How much do these off-diagonal artifacts affect the convergence speed?”. Figure 2 answers this question using the Rosenbrock function. Looking at the medians, the optimization reaches an objective value of $10^{-12}$ in $\sim 30$ iterations for Seq. Opt. ($B=1$), whereas C-BE with $B=2$ and $B\in\{5,10\}$ requires $\sim 50$ and more than $120$ iterations, respectively. Overall, MSO from $10$ initial points would require evaluations at about $300~(=30\times 10)$ points for $B=1$ and more than $1200=(120\times 10)$ for $B=10$, concluding that a naïve C-BE weakens the speedup effect. This is also the case for BFGS as shown in Figures 3–5 of Appendix B. It is worth noting that there is no principled way for L-BFGS-B to handle this phenomenon, although BFGS and L-BFGS have been extended to do so by griewank1982partitioned and bigeon2023framework, respectively. Namely, no work has addressed this issue for the acquisition function optimization, where each parameter has its bounds. The next section introduces a simple yet novel remedy that enables block-diagonal-structure-aware speedups without requiring new QN solvers.
Figure 2: Convergence speed of C-BE when using L-BFGS-B with the memory size $m=10$ on the Rosenbrock function ($D=5,\mathbf{x}\in[0,3]^{D}$). The figure shows the objective mean over $B$ restarts at each iteration. Each optimization is repeated $1000/B$ times. Each solid line and weak-color band represents the median and the $\pm$ IQR of the objective mean over $1000/B$ runs, respectively. Seq. Opt. corresponds to $B=1$. As the number of restarts $B$ increases, the convergence of C-BE requires substantially more iterations.
Algorithm 1 MSO: C-BE vs. D-BE
1:Inputs: $\alpha$ (Acquisition function), $B$ (number of restarts), $\mathbf{X}_{0}=\{\mathbf{x}^{(b)}_{0}\}_{b=1}^{B}$ (initial points)
2:Output: $\mathbf{x}^{\star}\in\mathop{\mathrm{argmax}}_{\mathbf{x}}\alpha(\mathbf{x})$
[…]
4:[C-BE] Initialize a QN optimizer $O_{\text{CBE}}$ on $\mathbf{X}\in\mathbb{R}^{B\times D}$
5:[D-BE] Initialize independent QN optimizers $O_{1},\dots,O_{B}$
6:while not converged do
7:  $\triangleright$ Batched Evaluation
8:  $\mathbf{a}\leftarrow\{\alpha(\mathbf{x}^{(b)})\}_{b=1}^{B}$, $\mathbf{g}\leftarrow\{\nabla\alpha(\mathbf{x}^{(b)})\}_{b=1}^{B}$
9:  $\triangleright$ QN Updates
10:  [C-BE]
11:  $\alpha_{\mathrm{sum}}\leftarrow\sum_{b=1}^{B}\alpha^{(b)}\coloneqq\sum_{b=1}^{B}\alpha(\mathbf{x}^{(b)})$
12:  $\mathbf{X}\leftarrow O_{\mathrm{CBE}}(\mathbf{X},\alpha_{\mathrm{sum}},\mathbf{g})$
13:  [D-BE]
14:  for $b=1,\dots,B$ do
15:  $\mathbf{x}^{(b)}\leftarrow O_{b}(\mathbf{x}^{(b)},\alpha^{(b)},\mathbf{g}^{(b)})$ return $\mathbf{x}^{\star}=\mathop{\mathrm{argmax}}_{\mathbf{x}\in\{\mathbf{x}^{(b)}\}_{b=1}^{B}}\alpha(\mathbf{x})$
4 Decoupling Updates Converge Faster without Difficult Mathematics
Why Is D-BE Better Than C-BE?
Algorithm 1 highlights the differences between D-BE (our proposal) and C-BE. D-BE is characterized by independent initialization and per-restart updates. Concretely, after each batched evaluation, D-BE performs $B$ independent L-BFGS-B updates—one per restart—using only its own history, thereby avoiding the off-diagonal artifacts present in C-BE.
Batching benefits not only acquisition-function evaluations but also the L-BFGS-B updates; however, the gain for the latter is limited. We now explain why. For example, a single evaluation of the expected improvement using a Gaussian process regressor with $n$ training data points costs $\mathcal{O}(n^{2}+nD)$, so evaluating $B$ points costs $\mathcal{O}(B(n^{2}+nD))$. By contrast, the L-BFGS-B update with the memory size $m$ for $B$ independent optimizations costs $\mathcal{O}(BmD)$. In regimes where $n\gg m$ (typically after tens to hundreds of trials, with $m\in[5,20]$), it is cost-effective to batch only evaluations while keeping each optimizer independent. As a result, D-BE avoids the off-diagonal artifacts induced by C-BE.
Most importantly, each restart in D-BE theoretically reproduces the per-restart update trajectory as in Seq. Opt. under the identical initialization and termination policies. Notice, however, that AD yields slightly different gradients between a batched evaluation and a single evaluation due to modulo floating-point nondeterminism, so the trajectory may not be identical. The general form of D-BE maintains a set of ongoing optimization indices $\mathcal{A}\subseteq\{1,\dots,B\}$, and prunes the converged optimizations. This mechanism allows D-BE to progressively shrink the batch size, reducing additional computational overhead. Note that Algorithm 1 does not detail the point-wise termination owing to brevity. In contrast, C-BE employs a QN optimizer with a shared QN state over all the restarts. Since detaching the converged optimizations is not feasible in C-BE, the overall runtime unnecessarily inflates.
Decouple L-BFGS-B Updates by Coroutine
We have discussed the benefit of using D-BE instead of C-BE. However, decoupling the L-BFGS-B updates often requires modifying the L-BFGS-B implementation directly. For example, SciPy provides L-BFGS-B while per-QN-iteration hooks are not exposed, making batched evaluations for independent optimizations non-trivial. Even in this scenario, we do not have to implement our own QN methods if we use a coroutine library. A coroutine is a special type of function that can pause its execution and be resumed later from the same point, allowing for cooperative multitasking and simplifying asynchronous programming. D-BE can be achieved by using a coroutine, which is composed of (1) one batch evaluator, and (2) $B$ workers for each independent optimization. In the coroutine, the evaluator performs a batched evaluation, it dispatches $(\alpha^{(b)},g^{(b)})$ to each worker, and each worker updates its QN state. Repeating this coroutine enables D-BE without any new solvers.
5 Benchmarking Experiments
We finally conduct the speedup effect and the solution quality check of D-BE against Seq. Opt. and C-BE through a BO application. The Rastrigin function from COCO (BBOB) suite (hansen2009real; hansen2021coco) available in OptunaHub (ozaki2025optunahub) is taken as an objective function for BO. The benchmarking results on other functions are available in Table 2 of Appendix. Note that the optimization is performed in the minimization direction in this section. Each BO iteration first fits a Gaussian process regressor with a Matérn-$\nu{=}5/2$ kernel on a set of observed pairs of a parameter vector and the corresponding objective value. Then, the next parameter vector is determined based on log expected improvement (LogEI). LogEI is optimized by MSO using L-BFGS-B in each BO iteration.
$D$ Method Best Value $\downarrow$ Runtime (s) $\downarrow$ Iters. $\downarrow$
5 Seq. Opt. $10.94$ $92.7$ $\mathbf{11.0}$
C-BE $11.28$ $62.0$ $35.0$
D-BE $\mathbf{10.85}$ $\mathbf{59.2}$ $\mathbf{11.0}$
10 Seq. Opt. $28.24$ $96.5$ $\mathbf{14.8}$
C-BE $24.71$ $73.7$ $54.2$
D-BE $\mathbf{23.89}$ $\mathbf{67.3}$ $15.0$
20 Seq. Opt. $59.08$ $129.4$ $21.5$
C-BE $57.04$ $95.4$ $68.8$
D-BE $\mathbf{55.81}$ $\mathbf{86.4}$ $\mathbf{20.0}$
40 Seq. Opt. $88.31$ $169.2$ $26.5$
C-BE $100.36$ $120.6$ $87.0$
D-BE $\mathbf{76.48}$ $\mathbf{108.2}$ $\mathbf{25.8}$
Table 1: The benchmarking results of BO with $300$ trials using L-BFGS-B with the memory size of $m=10$ and $B=10$ restarts on the $D$-dimensional Rastrigin function ($D=5,10,20,40$). The termination criterion of L-BFGS-B is $200$ iterations or $\|\nabla\alpha(\mathbf{x})\|_{\infty}\leq 10^{-2}$. Each column shows the median over 2020 independent runs, each with a different random seed. Best Value: the minimum objective value among the 300 trials minus the best objective value over all runs. Runtime (s): the total wall-clock time for BO. Iters.: the median of the L-BFGS-B iteration count over $300$ Trials $\times 10$ restarts. Lower is better.
Table 1 shows the benchmarking results. Based on the results, all methods achieved comparable (median) final objective values under the shared iteration cap. While C-BE occasionally outperformed Seq. Opt. for $D\in\{10,20\}$, we attribute this to seed variability and the median-of-minima aggregation rather than a systematic benefit of coupling, in contrast to the sizable iteration increase observed under C-BE. Meanwhile, as can be seen in the Iters. column, the L-BFGS-B iteration count drastically increased in C-BE. For example, C-BE required substantially more L-BFGS-B iterations than Seq. Opt. as $D$ grew, e.g., $68.8/21.5\!\approx\!3.2$ for $D=20$ and $87.0/26.5\!\approx\!3.3$ for D=40D{=}40. This observation is consistent with the discussion regarding the off-diagonal artifacts. These extra iterations can offset batching gains, making C-BE’s wall-clock time comparable to—and on some objectives worse than—Seq. Opt. By contrast, D-BE matches the iteration counts of the Seq. Opt. baseline while exploiting batched gradient evaluations, having reduced the runtime of the acquisition function optimization ($\sim\!1.5\times$ faster than Seq. Opt.). The same pattern holds on Sphere, Attractive Sector and Step Ellipsoidal from the COCO (BBOB) suite (hansen2009real; hansen2021coco); See Appendix C (Table 2) for more details.
6 Conclusion
This paper discussed the speedup effect of BO in light of the acquisition function optimization. Although the optimization of the summed acquisition function at multiple points has been known to be a fast implementation for MSO, we empirically showed that either the convergence speed or the solution quality is degraded owing to off-diagonal artifacts. In principle, C-BE degrades the solution quality when the iteration count is capped, or it takes significantly longer when we wait until the solution quality is satisfied. Mathematically speaking, a variant of L-BFGS-B can optimize efficiently as is, but such an extension is challenging. To this end, we proposed an explicit decoupling approach to circumvent the off-diagonal artifacts. In general, such decoupling requires L-BFGS-B to take a specific design especially when the BO implementation relies on PyTorch owing to its incompatibility with multi-processing. Meanwhile, our proposition only requires a coroutine to realize the decoupling. The experiments demonstrated that our approach accelerated BO while keeping the original performance.
Appendix
Appendix A Pseudocode: Sequential Optimization (Seq. Opt.)
Algorithm 2 MSO: Seq. Opt.
1:Inputs: $\alpha$ (Acquisition function), $B$ (number of restarts), $\mathbf{x}_{0}=\{\mathbf{x}^{(b)}_{0}\}_{b=1}^{B}$ (initial points)
2:Output: $\mathbf{x}^{\star}\in\mathop{\mathrm{argmax}}_{\mathbf{x}}\alpha(\mathbf{x})$
[…]
4:for $b=1,\dots,B$ do
5:  Initialize QN optimizer $O_{b}$; set $\mathbf{x}^{(b)}\leftarrow\mathbf{x}^{(b)}_{0}$
6:  while restart $b$ not converged do
7:   $\triangleright$ Evaluation (no batching across restarts)
8:   $a^{(b)}\leftarrow\alpha(\mathbf{x}^{(b)})$; $\mathbf{g}^{(b)}\leftarrow\nabla\alpha(\mathbf{x}^{(b)})$
9:   $\triangleright$ QN Updates
10:   $\mathbf{x}^{(b)}\leftarrow O_{b}\big(\mathbf{x}^{(b)},\,a^{(b)},\,\mathbf{g}^{(b)}\big)$   return $\mathbf{x}^{\star}=\mathop{\mathrm{argmax}}_{\{\mathbf{x}^{(b)}\}_{b=1}^{B}}\alpha(\mathbf{x})$
Algorithm 2 presents the Seq. Opt. algorithm, which serves as a baseline: each restart maintains its own optimizer state, and no batching is performed across restarts. As a result, no cross-restart information is shared, and the per-restart curvature is preserved by construction.
Appendix B BFGS Convergence on Separable Sums
We verify that the observed degradation under C-BE is not a limited-memory artifact: the same pattern appears with BFGS on Rosenbrock.
Figure 3: Same as Figure 1 but BFGS, i.e., memory is not limited, with B=3B{=}3, D=5D{=}5. (a) ground truth (block diagonal); (b) Seq. Opt.: structure preserved; (c) C-BE: dense off-diagonal couplings.
Figure 4: Same as Figure 1 but BFGS with B=10B{=}10, D=5D{=}5. Off-diagonal seems to be more serious with larger $B$.
Figure 5: Same as Figure 2 but BFGS. C-BE (B>1B{>}1) requires substantially more iterations than the Seq. Opt. baseline (B=1B{=}1 per restart).
We replicate the setup from Section 3 using BFGS with $B\in\{1,2,5,10\}$ restarts on D=5D{=}5 Rosenbrock function. Figures 34 show that C-BE again injects off-diagonal mass into the inverse Hessian approximation that should be zero for a separable sum objective, while Seq. Opt. preserves the block-diagonal structure. Consistently, iteration counts increase markedly under C-BE relative to Seq. Opt. (Figure 5).
These observations indicate that the failure arises from applying QN methods updates to a coupled yet separable problem (shared state across restarts), rather than from the limited-memory approximation.
Appendix C Full Benchmark Results
Objective ($D$) Method Best Value $\downarrow$ Runtime (s) $\downarrow$ Iters. $\downarrow$
Sphere (5) Seq. Opt. $4.026\times 10^{-5}$ $196.9$ $\mathbf{7.0}$
C-BE $\mathbf{1.068\times 10^{-5}}$ $\mathbf{121.8}$ $7.5$
D-BE $5.167\times 10^{-5}$ $128.2$ $\mathbf{7.0}$
Sphere (10) Seq. Opt. $9.632\times 10^{-4}$ $213.6$ $13.0$
C-BE $\mathbf{1.846\times 10^{-4}}$ $146.8$ $26.0$
D-BE $4.973\times 10^{-4}$ $\mathbf{141.6}$ $\mathbf{12.8}$
Sphere (20) Seq. Opt. $8.471\times 10^{-3}$ $225.4$ $15.5$
C-BE $9.820\times 10^{-3}$ $173.1$ $48.8$
D-BE $\mathbf{7.246\times 10^{-3}}$ $\mathbf{147.4}$ $\mathbf{15.0}$
Sphere (40) Seq. Opt. $0.04$ $291.4$ $\mathbf{21.2}$
C-BE $0.03$ $220.9$ $75.2$
D-BE $\mathbf{0.03}$ $\mathbf{190.2}$ $22.0$
AS (5) Seq. Opt. $32.53$ $127.8$ $\mathbf{12.0}$
C-BE $\mathbf{32.38}$ $\mathbf{72.2}$ $37.0$
D-BE $35.90$ $72.5$ $12.5$
AS (10) Seq. Opt. $\mathbf{59.69}$ $195.8$ $17.4$
C-BE $61.97$ $134.9$ $57.0$
D-BE $60.55$ $\mathbf{121.9}$ $\mathbf{17.0}$
AS (20) Seq. Opt. $55.75$ $203.8$ $\mathbf{6.2}$
C-BE $\mathbf{52.47}$ $183.2$ $104.0$
D-BE $61.16$ $\mathbf{151.6}$ $9.5$
AS (40) Seq. Opt. $319.00$ $130.5$ $\mathbf{3.0}$
C-BE $388.20$ $139.5$ $102.2$
D-BE $\mathbf{309.60}$ $\mathbf{108.1}$ $3.5$
SE (5) Seq. Opt. $0.09$ $115.5$ $\mathbf{12.5}$
C-BE $\mathbf{0.02}$ $70.6$ $37.0$
D-BE $0.10$ $\mathbf{69.4}$ $\mathbf{12.5}$
SE (10) Seq. Opt. $2.41$ $126.5$ $18.5$
C-BE $\mathbf{1.40}$ $89.2$ $75.0$
D-BE $2.37$ $\mathbf{77.4}$ $\mathbf{18.2}$
SE (20) Seq. Opt. $17.67$ $118.1$ $\mathbf{14.1}$
C-BE $21.78$ $95.2$ $76.0$
D-BE $\mathbf{17.56}$ $\mathbf{81.1}$ $14.8$
SE (40) Seq. Opt. $\mathbf{34.32}$ $293.9$ $\mathbf{23.0}$
C-BE $52.46$ $231.9$ $122.5$
D-BE $45.02$ $\mathbf{184.6}$ $\mathbf{23.0}$
Rastrigin (5) Seq. Opt. $10.94$ $92.7$ $\mathbf{11.0}$
C-BE $11.28$ $62.0$ $35.0$
D-BE $\mathbf{10.85}$ $\mathbf{59.2}$ $\mathbf{11.0}$
Rastrigin (10) Seq. Opt. $28.24$ $96.5$ $\mathbf{14.8}$
C-BE $24.71$ $73.7$ $54.2$
D-BE $\mathbf{23.89}$ $\mathbf{67.3}$ $15.0$
Rastrigin (20) Seq. Opt. $59.08$ $129.4$ $21.5$
C-BE $57.04$ $95.4$ $68.8$
D-BE $\mathbf{55.81}$ $\mathbf{86.4}$ $\mathbf{20.0}$
Rastrigin (40) Seq. Opt. $88.31$ $169.2$ $26.5$
C-BE $100.36$ $120.6$ $87.0$
D-BE $\mathbf{76.48}$ $\mathbf{108.2}$ $\mathbf{25.8}$
Table 2: Benchmarking results of BO ($300$ trials) using L-BFGS-B ($m=10$) with $B=10$ restarts on four objectives: Sphere, Attractive Sector (AS), Step Ellipsoidal (SE), and Rastrigin, each at $D\in\{5,10,20,40\}$. The Objective ($D$) column lists the objective name and dimensionality (e.g., Sphere (10) means $D=10$). The termination criterion of L-BFGS-B is $200$ iterations or $\|\nabla\alpha(\mathbf{x})\|_{\infty}\leq 10^{-2}$. Each entry is the median over 2020 independent runs with different seeds. Best Value: $\min$ objective among the $300$ trials minus the best objective over all runs. Runtime (s): total wall-clock time for BO. median L-BFGS-B iteration count over $300$ trials $\times$ $10$ restarts. Lower is better.
This appendix complements the main text (which reported only Rastrigin) by providing the complete results for Sphere, Attractive Sector, Step Ellipsoidal, and Rastrigin at $D\in\{5,10,20,40\}$ in Table 2. Experimental settings are identical to those stated in the caption and the main text.
Across objectives and dimensions, C-BE degrades notably in higher dimensions both in solution quality and in optimizer iterations. At D=40D{=}40, C-BE yields larger Best Value than D-BE by $25.4\%$ on Attractive Sector, $16.5\%$ on Step Ellipsoidal, and $31.3\%$ on Rastrigin (Sphere is roughly on par: $2.8\%$). [^2] Percentages computed as $(\textsc{C-BE}-\textsc{D-BE})/\textsc{D-BE}$ using Table 2. Its iteration counts also inflate sharply: at D=40D{=}40, C-BE requires about $3.3{\times}\!\sim\!29{\times}$ more L-BFGS-B iterations than D-BE ($3.3{\times}\!\sim\!34{\times}$ vs. Seq. Opt.), with the largest gap on Attractive Sector (102.2 vs. 3.5 and 3.0).
[…]

### v2

@@
 Our approach not only yields the theoretically identical convergence to the sequential MSO but also drastically reduces the wall-clock time compared to the previous approaches.
-Codehttps://github.com/Kaichi-Irie/aaai2026-anon-exp/tree/submission
+Our approach is available in GPSampler in Optuna, effectively reducing its computational overhead.
+Codehttps://github.com/Kaichi-Irie/faster-batched-acqf-opt-experiments/tree/submission
 1 Introduction
@@
     Demonstrate that D-BE is faster than Seq. Opt. and C-BE in experiments.
+Notice that our approach has been successfully merged into GPSampler in Optuna, significantly accelerating its runtime.
 Figure 1: Contour maps of the inverse Hessian (Left) and the inverse Hessians approximated by L-BFGS-B with Seq.
@@
 Based on the results, all methods achieved comparable (median) final objective values under the shared iteration cap.
-While C-BE occasionally outperformed Seq.
-Opt.
-for $D\in\{10,20\}$, we attribute this to seed variability and the median-of-minima aggregation rather than a systematic benefit of coupling, in contrast to the sizable iteration increase observed under C-BE.
-Meanwhile, as can be seen in the Iters.
+As can be seen in the Iters.
 column, the L-BFGS-B iteration count drastically increased in C-BE.
@@
 Opt.).
-The same pattern holds on Sphere, Attractive Sector and Step Ellipsoidal from the COCO (BBOB) suite (hansen2009real; hansen2021coco); See Appendix C (Table 2) for more details.
+The similar pattern holds on Sphere, Attractive Sector and Step Ellipsoidal from the COCO (BBOB) suite (hansen2009real; hansen2021coco); See Appendix C (Table 2) for more details.
+Note that while C-BE occasionally outperformed Seq.
+Opt.
+for $D\in\{10,20\}$, we attribute this to the difference in the convergence behavior rather than a systematic benefit of coupling.
+More specifically, the acquisition function optimization for C-BE is mostly terminated by the gradient tolerance but not the iteration count, meaning that the optimization results for C-BE are not premature.
+This leads to very different optimization results between C-BE and Seq.
+Opt., making it possible for C-BE to sometimes outperform Seq.
+Opt.
 6 Conclusion
@@
 Appendix
-Appendix A Pseudocode: Sequential Optimization (Seq. Opt.)
+Appendix A Sequential Optimization (Seq. Opt.)
 Algorithm 2 MSO: Seq. Opt.
@@
 Opt.
-algorithm, which serves as a baseline: each restart maintains its own optimizer state, and no batching is performed across restarts.
+algorithm, which serves as a baseline.
+Each restart maintains its own optimizer state, and no batching is performed across restarts.
 As a result, no cross-restart information is shared, and the per-restart curvature is preserved by construction.
-Appendix B BFGS Convergence on Separable Sums
-We verify that the observed degradation under C-BE is not a limited-memory artifact: the same pattern appears with BFGS on Rosenbrock.
-Figure 3: Same as Figure 1 but BFGS, i.e., memory is not limited, with B=3B{=}3, D=5D{=}5.
-(a) ground truth (block diagonal); (b) Seq.
-Opt.: structure preserved; (c) C-BE: dense off-diagonal couplings.
-Figure 4: Same as Figure 1 but BFGS with B=10B{=}10, D=5D{=}5. Off-diagonal seems to be more serious with larger $B$.
-Figure 5: Same as Figure 2 but BFGS. C-BE (B>1B{>}1) requires substantially more iterations than the Seq. Opt. baseline (B=1B{=}1 per restart).
-We replicate the setup from Section 3 using BFGS with $B\in\{1,2,5,10\}$ restarts on D=5D{=}5 Rosenbrock function.
-Figures 34 show that C-BE again injects off-diagonal mass into the inverse Hessian approximation that should be zero for a separable sum objective, while Seq.
+Appendix B Off-Diagonal Artifact Effect on BFGS
+Figure 3: Contour maps of the inverse Hessian (Left) and its approximations by Seq.
+Opt.
+(Center) and C-BE (Right).
+This figure follows the same setup as Figure 1 except that BFGS, i.e., the memory is not limited, is used instead of L-BFGS-B.
+Figure 4: Contour maps of the inverse Hessian (Left) and its approximations by Seq.
+Opt.
+(Center) and C-BE (Right).
+This figure follows the same setup as Figure 1 except that BFGS, i.e., the memory is not limited, is used instead of L-BFGS-B, and $B=10$ is used instead of $B=3$.
+Off-diagonal artifacts are more prominent for a large $B$.
+Figure 5: Convergence speed of C-BE when using BFGS.
+This figure follows the same setup as Figure 2 except that BFGS is used instead of L-BFGS-B.
+As the number of restarts $B$ increases, the convergence of C-BE requires substantially more iterations.
+This section verifies that the observed degradation under C-BE does not stem from limiting the memory of BFGS.
+The setup in this section follows Section 3.
+Figures 3 and 4 show that C-BE injects off-diagonal mass into the inverse Hessian approximation, which should be zero, while Seq.
 Opt.
@@
 Opt.
-(Figure 5).
-These observations indicate that the failure arises from applying QN methods updates to a coupled yet separable problem (shared state across restarts), rather than from the limited-memory approximation.
-Appendix C Full Benchmark Results
-Objective ($D$) Method Best Value $\downarrow$ Runtime (s) $\downarrow$ Iters. $\downarrow$
-Sphere (5) Seq. Opt. $4.026\times 10^{-5}$ $196.9$ $\mathbf{7.0}$
+as seen in Figure 5.
+These observations confirm that the off-diagonal artifacts are caused by coupling QN updates for a separable problem rather than by limiting the memory.
+Appendix C Full Benchmarking Results
+Objective Method Best Value $\downarrow$ Runtime (s) $\downarrow$ Iters. $\downarrow$
+5DD Sphere Seq. Opt. $4.026\times 10^{-5}$ $196.9$ $\mathbf{7.0}$
 C-BE $\mathbf{1.068\times 10^{-5}}$ $\mathbf{121.8}$ $7.5$
 D-BE $5.167\times 10^{-5}$ $128.2$ $\mathbf{7.0}$
-Sphere (10) Seq. Opt. $9.632\times 10^{-4}$ $213.6$ $13.0$
+10DD Sphere Seq. Opt. $9.632\times 10^{-4}$ $213.6$ $13.0$
 C-BE $\mathbf{1.846\times 10^{-4}}$ $146.8$ $26.0$
 D-BE $4.973\times 10^{-4}$ $\mathbf{141.6}$ $\mathbf{12.8}$
-Sphere (20) Seq. Opt. $8.471\times 10^{-3}$ $225.4$ $15.5$
+20DD Sphere Seq. Opt. $8.471\times 10^{-3}$ $225.4$ $15.5$
 C-BE $9.820\times 10^{-3}$ $173.1$ $48.8$
 D-BE $\mathbf{7.246\times 10^{-3}}$ $\mathbf{147.4}$ $\mathbf{15.0}$
-Sphere (40) Seq. Opt. $0.04$ $291.4$ $\mathbf{21.2}$
+40DD Sphere Seq. Opt. $0.04$ $291.4$ $\mathbf{21.2}$
 C-BE $0.03$ $220.9$ $75.2$
 D-BE $\mathbf{0.03}$ $\mathbf{190.2}$ $22.0$
-AS (5) Seq. Opt. $32.53$ $127.8$ $\mathbf{12.0}$
+5DD AS Seq. Opt. $32.53$ $127.8$ $\mathbf{12.0}$
 C-BE $\mathbf{32.38}$ $\mathbf{72.2}$ $37.0$
 D-BE $35.90$ $72.5$ $12.5$
-AS (10) Seq. Opt. $\mathbf{59.69}$ $195.8$ $17.4$
+10DD AS Seq. Opt. $\mathbf{59.69}$ $195.8$ $17.4$
 C-BE $61.97$ $134.9$ $57.0$
 D-BE $60.55$ $\mathbf{121.9}$ $\mathbf{17.0}$
-AS (20) Seq. Opt. $55.75$ $203.8$ $\mathbf{6.2}$
+20DD AS Seq. Opt. $55.75$ $203.8$ $\mathbf{6.2}$
 C-BE $\mathbf{52.47}$ $183.2$ $104.0$
 D-BE $61.16$ $\mathbf{151.6}$ $9.5$
-AS (40) Seq. Opt. $319.00$ $130.5$ $\mathbf{3.0}$
+40DD AS Seq. Opt. $319.00$ $130.5$ $\mathbf{3.0}$
 C-BE $388.20$ $139.5$ $102.2$
 D-BE $\mathbf{309.60}$ $\mathbf{108.1}$ $3.5$
-SE (5) Seq. Opt. $0.09$ $115.5$ $\mathbf{12.5}$
+5DD SE Seq. Opt. $0.09$ $115.5$ $\mathbf{12.5}$
 C-BE $\mathbf{0.02}$ $70.6$ $37.0$
 D-BE $0.10$ $\mathbf{69.4}$ $\mathbf{12.5}$
-SE (10) Seq. Opt. $2.41$ $126.5$ $18.5$
+10DD SE Seq. Opt. $2.41$ $126.5$ $18.5$
 C-BE $\mathbf{1.40}$ $89.2$ $75.0$
 D-BE $2.37$ $\mathbf{77.4}$ $\mathbf{18.2}$
-SE (20) Seq. Opt. $17.67$ $118.1$ $\mathbf{14.1}$
+20DD SE Seq. Opt. $17.67$ $118.1$ $\mathbf{14.1}$
 C-BE $21.78$ $95.2$ $76.0$
 D-BE $\mathbf{17.56}$ $\mathbf{81.1}$ $14.8$
-SE (40) Seq. Opt. $\mathbf{34.32}$ $293.9$ $\mathbf{23.0}$
+40DD SE Seq. Opt. $\mathbf{34.32}$ $293.9$ $\mathbf{23.0}$
 C-BE $52.46$ $231.9$ $122.5$
 D-BE $45.02$ $\mathbf{184.6}$ $\mathbf{23.0}$
-Rastrigin (5) Seq. Opt. $10.94$ $92.7$ $\mathbf{11.0}$
+5DD Rastrigin Seq. Opt. $10.94$ $92.7$ $\mathbf{11.0}$
 C-BE $11.28$ $62.0$ $35.0$
 D-BE $\mathbf{10.85}$ $\mathbf{59.2}$ $\mathbf{11.0}$
-Rastrigin (10) Seq. Opt. $28.24$ $96.5$ $\mathbf{14.8}$
+10DD Rastrigin Seq. Opt. $28.24$ $96.5$ $\mathbf{14.8}$
 C-BE $24.71$ $73.7$ $54.2$
 D-BE $\mathbf{23.89}$ $\mathbf{67.3}$ $15.0$
-Rastrigin (20) Seq. Opt. $59.08$ $129.4$ $21.5$
+20DD Rastrigin Seq. Opt. $59.08$ $129.4$ $21.5$
 C-BE $57.04$ $95.4$ $68.8$
 D-BE $\mathbf{55.81}$ $\mathbf{86.4}$ $\mathbf{20.0}$
-Rastrigin (40) Seq. Opt. $88.31$ $169.2$ $26.5$
+40DD Rastrigin Seq. Opt. $88.31$ $169.2$ $26.5$
 C-BE $100.36$ $120.6$ $87.0$
 D-BE $\mathbf{76.48}$ $\mathbf{108.2}$ $\mathbf{25.8}$
-Table 2: Benchmarking results of BO ($300$ trials) using L-BFGS-B ($m=10$) with $B=10$ restarts on four objectives: Sphere, Attractive Sector (AS), Step Ellipsoidal (SE), and Rastrigin, each at $D\in\{5,10,20,40\}$.
-The Objective ($D$) column lists the objective name and dimensionality (e.g., Sphere (10) means $D=10$).
+Table 2: The benchmarking results of BO with $300$ trials using L-BFGS-B with the memory size of $m=10$ and $B=10$ restarts on the $D$-dimensional functions ($D=5,10,20,40$).
+We picked four objectives: Sphere, Attractive Sector (AS), Step Ellipsoidal (SE), and Rastrigin.
 The termination criterion of L-BFGS-B is $200$ iterations or $\|\nabla\alpha(\mathbf{x})\|_{\infty}\leq 10^{-2}$.
-Each entry is the median over 2020 independent runs with different seeds.
-Best Value: $\min$ objective among the $300$ trials minus the best objective over all runs.
-Runtime (s): total wall-clock time for BO.
-median L-BFGS-B iteration count over $300$ trials $\times$ $10$ restarts.
+Each column shows the median over 2020 independent runs, each with a different random seed.
+Best Value: the minimum objective value among the 300 trials minus the best objective value over all runs.
+Runtime (s): the total wall-clock time for BO.
+Iters.: the median of the L-BFGS-B iteration count over $300$ Trials $\times 10$ restarts.
 Lower is better.
-This appendix complements the main text (which reported only Rastrigin) by providing the complete results for Sphere, Attractive Sector, Step Ellipsoidal, and Rastrigin at $D\in\{5,10,20,40\}$ in Table 2.
-Experimental settings are identical to those stated in the caption and the main text.
-Across objectives and dimensions, C-BE degrades notably in higher dimensions both in solution quality and in optimizer iterations.
-At D=40D{=}40, C-BE yields larger Best Value than D-BE by $25.4\%$ on Attractive Sector, $16.5\%$ on Step Ellipsoidal, and $31.3\%$ on Rastrigin (Sphere is roughly on par: $2.8\%$).
-[^2] Percentages computed as $(\textsc{C-BE}-\textsc{D-BE})/\textsc{D-BE}$ using Table 2.
-Its iteration counts also inflate sharply: at D=40D{=}40, C-BE requires about $3.3{\times}\!\sim\!29{\times}$ more L-BFGS-B iterations than D-BE ($3.3{\times}\!\sim\!34{\times}$ vs.
-Seq.
-Opt.), with the largest gap on Attractive Sector (102.2 vs.
-3.5 and 3.0).
+Table 2 provides the complete results for Sphere, Attractive Sector, Step Ellipsoidal, and Rastrigin.
+Note that the setup follows Section 5.
+The results show that C-BE degrades notably in higher dimensions both in solution quality and in optimizer iterations.
+For example, C-BE yields larger Best Value in 40DD than D-BE by $25.4\%$ on Attractive Sector, $16.5\%$ on Step Ellipsoidal, $31.3\%$ on Rastrigin, and $2.8\%$ on Sphere  [^2] Percentages computed as $(\textsc{C-BE}-\textsc{D-BE})/\textsc{D-BE}$ using Table 2.
+.
+The iteration counts of C-BE also inflate sharply at D=40D{=}40, requiring about $3.3{\times}\!\sim\!29{\times}$ more L-BFGS-B iterations than D-BE.
+The largest gap is observed on Attractive Sector (102.2 vs.
+3.5).
 […]